# Enhancing the Robustness of LLM-Generated Code: Empirical Study and Framework

### v3

@@ Abstract.
Ensuring the robustness of code generated by large language models (LLMs) is crucial for real-world reliability. However, existing evaluations predominantly focus on correctness, often neglecting key robustness concerns such as missing input validation and insufficient error handling.
In this paper, we present the first empirical study on the robustness of LLM-generated code using the CoderEval benchmark. We introduce novel robustness metrics and analyze four state-of-the-art code LLMs, revealing that, on average, 43.1% of their generated code is less robust than human-written counterparts. Notably, over 90% of robustness deficiencies stem from missing conditional checks, with 70% of these omissions occurring in the first line of code. Additionally, in 69% of cases where a conditional statement is necessary but absent, the “if” token still ranks third or higher in the model’s predicted token probabilities, indicating an implicit recognition of control structures.
Building on these findings, we propose RobGen, a framework designed to enhance code robustness without requiring model retraining. RobGen leverages two model-agnostic techniques: RobGen-Adj, which dynamically adjusts token probabilities during decoding to encourage the inclusion of control structures, and RobGen-Ins, which improves generated code by inserting missing conditionals after generation. Experimental results demonstrate that RobGen reduces the proportion of less robust model-generated code by 20.0%, significantly enhancing code reliability across diverse tasks. As a lightweight and adaptable solution, RobGen effectively mitigates robustness challenges in LLM-generated code. All code and data are available at https://github.com/SYSUSELab/RobGen.
@@ 1. Introduction
Empirical Study. This work presents the first comprehensive study on the robustness of LLM-generated code using CoderEval, a benchmark for complex repository-level code generation. New robustness metrics are designed to assess code properties directly, moving beyond reliance on unit tests by incorporating comparisons with human-written ground truth. The study analyzes four Code LLMs, investigating (1) the robustness gap between generated and human-written code, (2) prevalent patterns of robustness issues, (3) the specific locations where these issues arise, and (4) whether LLMs inherently recognize the need for robustness checks.
Our empirical study reveals that on average 43.1% of LLM-generated code is less robust than human-written code, with significant room for improvement. We identify nine distinct patterns of robustness issues, 90% of which are related to missing conditional checks. Further analysis shows that 70% of these issues occur in the first line of the generated code. Additionally, in these patterns, the ”if” token ranks third or higher in 69% of the cases where conditionals should be generated but are missing.
Plug-in Framework. Building on our findings, we introduce RobGen, a plug-in framework for enhancing the robustness of LLM-generated code without retraining. RobGen consists of two model-agnostic independent techniques: RobGen-Adj and RobGen-Ins. RobGen-Adj operates during decoding, dynamically adjusting token logits to encourage the generation of essential control structures. RobGen-Ins refines code post-generation by identifying and inserting missing condition checks. As a lightweight and adaptable solution, RobGen effectively improves code reliability across different models and tasks while fully leveraging the model’s inherent capabilities. We propose the Robustness Relative Index (RRI) to measure the robustness difference between generated and reference code. An RRI below zero (RRI ¡ 0) means the model-generated code is less robust. Experimental results across five models show that RobGen reduces the proportion of code with RRI ¡ 0 by 20%, demonstrating its effectiveness.
@@
  • Proposing quantitative metrics to assess robustness, focusing on edge cases, invalid inputs, and error handling.
  • Introducing RobGen, a plug-in framework with RobGen-Adj and RobGen-Ins, two lightweight, model-agnostic techniques that refine control structures during and after generation without retraining.
@@ 2. Empirical Study
  • RQ1: How robust is the code generated by LLMs compared to human-written code? We evaluate LLM-generated code robustness by comparing it to human-written code, focusing on handling boundary conditions, invalid inputs, and unexpected runtime scenarios.
@@ 2.1.3. Implementation Details
2.2. RQ1: Robustness of LLM-Generated vs. Human Code
To answer this RQ, we compare the robustness of LLM-generated code with the ground truth implementations, which are human-written code from real-world projects.
@@ 2.2.1. Analysis Target Selection
As described in Section 2.1, we use 230 Java coding tasks from CoderEval and generate one code snippet per task using different models with greedy decoding. Table 1 presents the Compile@1 (i.e., the number of generated code snippets that successfully compile without errors) and Pass@1 (i.e., the number of snippets that pass all test cases), along with the corresponding counts for each model.
We first filter out LLM-generated code that fail to compile as such code is often incomplete or contains undefined methods and variables. However, our evaluation extends beyond snippets that pass all test cases. We observe that some code failures result not from logic errors but from missing robustness checks, such as improper handling of boundary conditions (e.g., A <= B vs. A < B). These cases, which fail tests due to a lack of robustness, are central to our RQ1 analysis and are crucial for assessing the robustness of model-generated code.
@@
Qwen2.5-Coder-1.5B 0.67 (155) 0.39 (90)
Qwen2.5-Coder-7B 0.74 (171) 0.49 (112)
2.2.2. Analysis Metrics
To evaluate the robustness of generated code, we adopt principles from defensive programming (McConnell, 2004), which emphasizes boundary checking, exception handling, and input validation to ensure reliable software behavior. Robust code anticipates and mitigates potential errors by systematically handling abnormal conditions. Our analysis focuses on two key aspects: control expressions and exception handling.
Control Expressions Analysis. To evaluate the structure and complexity of control logic, we analyze atomic Boolean expressions, which are the smallest evaluable Boolean expressions found in conditional statements (e.g., if, for, while). These include simple conditions such as x > 5 or isValid() that serve as fundamental decision-making units in code. Using abstract syntax trees (ASTs) extracted via tree-sitter (Hendrik van Antwerpen, 2022), we systematically identify and extract all atomic Boolean expressions from the generated and reference code. This enables us to quantitatively assess the robustness of control structures.
Atomic Boolean Expression Extraction Method. When extracting atomic Boolean expressions, if the expression involves local variables, we replace them with their corresponding values or expressions in the code. Specifically, if a local variable is assigned a value from a method call or an operation, the atomic Boolean expression will be rewritten to reflect that operation directly, avoiding any reference to intermediate variables. For example, consider the code int len = a.length; and if (len > 0). In this case, the atomic Boolean expression extracted from the if statement would be a.length > 0 instead of len > 0. This ensures that the extracted atomic Boolean expressions directly reflect the logic being tested in the code, without unnecessary intermediate variables, making the analysis more accurate.
The following metrics are defined to facilitate this analysis:
Average Atomic Boolean Expressions (AvgABE): The AvgABE measures the average number of atomic Boolean expressions per code snippet. It is calculated as:
@@ 2.2.2. Analysis Metrics
where $U_{i}$ represents the set of atomic Boolean expressions in code snippet $i$, and $N$ is the total number of snippets. A higher AvgABE value suggests that the code contains more explicit conditions, potentially indicating stronger robustness through enhanced boundary checking and input validation.
Atomic Boolean Expression Similarity (ABES): We define ABES to measure the proportion of shared atomic Boolean expressions between the generated code and the reference code. Given the sets of atomic Boolean expressions in the generated code ($U_{A}$) and the reference code ($U_{B}$), ABES is computed as:
(2) $ABES(A,B)=\frac{|U_{A}\cap U_{B}|}{|U_{B}|}$
When calculating the intersection between $U_{A}$ and $U_{B}$, we account for potential differences in how semantically equivalent atomic Boolean expressions are written. For instance, expressions like len > 0 and len $\geq$ 1, or File.exists() == True and File.exists() $\neq$ False, while syntactically different, may express equivalent logic. To handle these cases, we apply a set of predefined rules to perform approximate matching between these expressions. This metric quantifies the extent to which the generated code aligns with human-written code in terms of control expression structure. A higher ABES value suggests that the generated code adopts similar robustness practices, such as boundary checks and validation logic, as seen in human-written code.
Relative Robustness Index (RRI): The RRI quantifies the relative robustness of generated code $A$ compared to reference code $B$ by analyzing atomic Boolean expressions in control structures. This metric not only measures the extent to which the generated code preserves the robustness mechanisms of the reference but also evaluates whether it introduces additional safeguards. A higher RRI indicates that the generated code enhances robustness by incorporating new checks, while a lower value suggests a loss of essential robustness features.
As defined in Eq 3,
(3) $RRI(A,B)=\frac{|U_{A}\cap U_{B}|+\alpha|U_{\text{extra}}|}{|U_{B}|}-1$
where $U_{A}$ and $U_{B}$ represent the sets of atomic Boolean expressions in the generated and reference code, respectively, and $U_{\text{extra}}=U_{A}\setminus U_{B}$ captures additional robustness checks unique to the generated code. The weighting factor $\alpha$ (set to 1) modulates the impact of these extra expressions. An RRI of 0 indicates that the generated and reference code exhibit comparable robustness. A positive RRI suggests that the generated code strengthens robustness beyond the reference, while a negative value implies that the generated code omits some robustness measures present in the reference. By jointly considering both retained and newly introduced robustness checks, RRI provides a structured and interpretable measure for evaluating robustness improvements or degradations in generated code.
Exception Handling Analysis Exception handling is a key indicator of robustness. Our analysis focuses on two aspects: the adoption rate of exception handling and its consistency with the reference code.
Exception Handling Adoption Rate (EHAR): EHAR quantifies the proportion of code snippets that include try-catch blocks:
@@
A higher EHAR suggests a greater reliance on exception handling, which can be beneficial for capturing runtime errors but may indicate excessive dependence on reactive mechanisms rather than preventive checks.
Exception Handling Consistency (EHC): EHC measures the alignment of exception handling strategies between generated and reference code:
(5) $EHC(A,B)=\frac{|EH_{A}\cap EH_{B}|}{|EH_{B}|}$
where $EH_{A}$ and $EH_{B}$ represent the sets of snippets using try-catch in the generated and reference code, respectively. A lower EHC value suggests a deviation from human-written exception handling patterns, indicating either overuse or underuse of structured error handling.
Table 2. Comparison of Robustness Metrics Across Models: “DSC” for DeepSeekCoder and “QWC” for Qwen2.5-Coder
Model EHAR EHC AvgABE ABSE
@@
DSC-1.3B 0.02 0.03 0.40 1.39 2.04 0.42
DSC-6.7B 0.03 0.02 0.75 1.40 2.10 0.46
QWC-1.5B 0.01 0.03 0.25 1.62 2.10 0.48
QWC-7B 0.03 0.03 0.80 1.78 2.02 0.48
Figure 2. RRI of Different LLMs. RRI ¡ 0: Less robust; RRI ¿ 0: More robust (vs. reference).
@@ 2.2.3. Results
For the selected LLM-generated code (that can compile), we computed control expression and exception handling metrics relative to the reference code. The results are presented in Table 2 and Figure 2.
Control Expression Analysis. The AvgABE metric quantifies the average number of atomic Boolean expressions per code snippet. As indicated in Table 2, for the ground truth code, the AvgABE values across different subsets of model-generated code consistently remain around 2.0, suggesting that each ground truth snippet contains approximately two atomic Boolean expressions on average. In contrast, model-generated code tends to exhibit lower AvgABE values. Notably, DeepSeekCoder-1.3B demonstrates the lowest AvgABE value of 1.39. Among the evaluated models, Qwen2.5-Coder-7B performs relatively well, achieving an AvgABE of 1.78.
The ABES metric measures the proportion of atomic Boolean expressions shared between the generated and reference code, relative to the total number of atomic Boolean expressions in the reference code. The results indicate that, across different models, ABES values are relatively consistent, suggesting that despite variations in the average number of atomic Boolean expressions across models, the overlap between generated and reference code in this aspect remains stable. The only exception is DeepSeekCoder-1.3B, which exhibits a significantly lower ABES value of 0.42. For other models, the overlap remains around 50%.
Figure 2 shows the distribution of RRI ¡ 0, RRI = 0, and RRI ¿ 0 across models and tasks, representing code with lower robustness, equal robustness, and better robustness compared to the reference code, respectively. Overall, 33.9%-52.3% (43.1% on average) of the code is RRI¡0, 36.0%-49.7% (42.9% on average) has an RRI of 0, and 11.8%-16.4% (14.1% on average) has an RRI ¿ 0. The gap between the proportions of RRI ¡ 0 and RRI ¿ 0 highlights significant potential for improvement.
For Qwen2.5-Coder-1.5B, 45.8% of compiled code is less robust than the reference code, while for Qwen2.5-Coder-7B, this proportion decreases to 33.9%. This suggests that increasing the parameter size of the Qwen2.5-Coder models leads to improved robustness in generated code. These results suggest that, overall, larger models tend to generate more robust code. When comparing models of similar parameter sizes, Qwen2.5-Coder exhibits superior robustness in code generation compared to DeepSeekCoder. Specifically, for Qwen2.5-Coder-1.5B, the proportion of code with RRI = 0 is 38.7%, and the proportion with RRI ¿ 0 is 15.5%. In contrast, for DeepSeekCoder-1.3B, these proportions are 36.0% and 11.8%, respectively, indicating slightly weaker robustness.
@@
The EHC metric measures the consistency of exception handling between generated and reference code. Qwen2.5-Coder-7B scores 0.8, indicating strong alignment with human-authored patterns, while Qwen2.5-Coder-1.5B scores 0.25, suggesting difficulty in replicating ground truth exception handling.
{mdframed}
[linecolor=gray!25,roundcorner=12pt,backgroundcolor=myyellow!30,linewidth=3pt,innerleftmargin=2pt, leftmargin=0cm,rightmargin=0cm,topline=false,bottomline=false,rightline = false] Finding 1: Analysis of robustness-related metrics shows that, compared to human-written code, LLM-generated code is suboptimal, with an average of 43.1% being less robust. There is significant potential for improvement. Among various factors, control expressions require more focus than exception handling, as they are more common in real-world code and exhibit a larger performance gap in LLM-generated code.
@@ 2.3.1. Data Selection
The dataset used for this analysis originates from RQ1, consisting of 920 generated code snippets from four models (230 snippets per model). To efficiently identify robustness issues, we leverage the Relative Robustness Index (RRI) from RQ1 2.2.2, selecting only snippets where $RRI<0$ for further analysis. Given the high cost of full manual analysis, we use RRI as an initial filtering step to retain only code snippets exhibiting lower robustness than the reference implementation, resulting in a refined dataset with 437 code snippets for detailed manual inspection.
Further, we manually examine each retained snippet and filter out cases where the generated code is entirely incorrect and incomparable to the reference. For example, a snippet with compilation errors due to formatting inconsistencies in the return value but otherwise correct logic is kept, while one that fails to implement the required functionality is discarded. As a result, 400 code snippets are kept for further analysis.

### v4

# A Preliminary Study on the Robustness of Code Generation by Large Language Models

@@ Abstract.
Robustness is a critical factor for reliable code generation by large language models, yet most evaluations focus on correctness and overlook key issues such as missing input validation and inadequate error handling. In this work, we present the first empirical study of LLM-generated code robustness using the CoderEval benchmark. Evaluating four state-of-the-art code LLMs, we find that 35.2% of their outputs are less robust than human-written code, with over 90% of deficiencies caused by missing conditional checks—70% of which occur in the first line. Interestingly, in 63% of cases where a conditional statement is needed but absent, the “if” token still ranks among the top three predictions, suggesting implicit recognition of control flow.
To address these issues, we propose RobGen, a model-agnostic framework that improves robustness without retraining. RobGen combines a line-level intervention checker, which decides whether to adjust logits for each generated line, with token-level conditional logit adjustments to promote essential control structures. Experiments show that RobGen reduces the proportion of less robust code by 10%, achieves the highest average Pass@1 (43.57), and adds minimal overhead (+33.4%). As a lightweight and adaptable solution, RobGen effectively enhances the reliability of LLM-generated code across diverse tasks.
@@ 1. INTRODUCTION
Empirical Study.This work presents the first comprehensive study on the robustness of LLM-generated code using CoderEval, a benchmark for complex repository-level code generation. Unlike prior approaches that primarily rely on unit tests, our methodological contribution lies in leveraging LLM as evaluator to directly assess code properties, complemented by comparisons with human-written ground truth for a more comprehensive and fine-grained robustness evaluation. The study analyzes four Code LLMs, investigating (1) the robustness gap between generated and human-written code, (2) prevalent patterns of robustness issues, (3) the specific locations where these issues arise, and (4) whether LLMs inherently recognize the need for robustness checks.
Our empirical study reveals that on average 37.8% of LLM-generated code is less robust than human-written code, with significant room for improvement. We identify nine distinct patterns of robustness issues, 90% of which are related to missing conditional checks. Further analysis shows that 70% of these issues occur in the first line of the generated code. Additionally, in these patterns, the ”if” token ranks third or higher in 69% of the cases where conditionals should be generated but are missing.
Plug-in Framework. Building on our findings, we introduce RobGen, a plug-and-play framework for enhancing the robustness of LLM-generated code without retraining. RobGen consists of two model-agnostic techniques: line-level intervention checker and token-level conditional logit adjustment. During code generation, the line-level intervention checker inspects each line to decide if a logit adjustment should be applied. token-level conditional logit adjustment operates during decoding, dynamically adjusting token logits to encourage the generation of essential control structures. As a lightweight and adaptable solution, RobGen effectively improves code reliability across different models and tasks while fully leveraging the model’s inherent capabilities. We leverage LLM as evaluator to measure the robustness difference between generated and reference code. Experimental results across five models show that RobGen reduces the proportion of less robust code by 10% and achieves the highest average Pass@1 (43.57), enhancing code generation robustness with minimal time overhead (+33.4%), thereby demonstrating its effectiveness.
@@
  • Leveraging LLM to assess robustness, focusing on edge cases, invalid inputs, and error handling.
  • Introducing RobGen, a plug-in framework with two lightweight, model-agnostic techniques that refine control structures during and after generation without retraining.
@@ 2. EMPIRICAL STUDY
  • RQ1: How robust is the code generated by LLMs compared to human-written code? We employ an LLM-based evaluator to assess the robustness of LLM-generated code by comparing it to human-written code, and further analyze statistical metrics to capture concrete aspects of robustness.
@@ 2.1.3. Implementation Details
2.2. RQ1: Robustness of LLM-Generated Vs. Human-Written Code
To answer this RQ, we compare the robustness of LLM-generated code with the ground truth implementations, which are human-written code from real-world projects. For the robustness evaluation, we employ both a LLM as an automatic evaluator and complementary statistical metrics to provide quantitative evidence.
@@ 2.2.1. Analysis Target Selection
As described in Section 2.1, we use 230 Java coding tasks from CoderEval and generate one code snippet per task using different models with greedy decoding. Table 1 presents the Compile@1 (i.e., the number of generated code snippets that successfully compile without errors) and Pass@1 (i.e., the number of snippets that pass all test cases), along with the corresponding counts for each model. We filter out LLM-generated code that fail to compile as such code is often incomplete or contains undefined methods and variables.
2.2.2. LLM-Based Automatic Evaluator
@@
Qwen2.5-Coder-1.5B 0.67 (154) 0.40 (91)
Qwen2.5-Coder-7B 0.75 (172) 0.49 (112)
To evaluate code robustness at scale, we adopt a LLM as an automatic evaluator. Using LLMs as “judges” has become common in tasks such as code quality assessment, summarization evaluation, and natural language generation benchmarking (Busker et al., 2025; Sufi, 2024), due to their strong analytical and contextual reasoning abilities. Compared with manual evaluation, this approach is cost-efficient, reproducible, and avoids inter-rater variability (Cheng et al., 2023).
We employ GPT-4o as the evaluator to compare two code snippets (LLM-generated vs. human-written). Fig. 2 illustrates the prompt template for assessing a code pair in a specified language. The evaluation follows defensive programming principles (McConnell, 2004), which emphasize boundary checking, exception handling, and input validation—core practices that directly relate to robustness. Guided by these criteria, the prompt instructs the LLM to assign a comparative robustness score. To avoid position bias, Code A and Code B are randomly assigned, hiding their origin from the evaluator. Scores are on a 1–5 scale: >3>3 indicates Code A is more robust, <3<3 less robust, and 3 denotes comparable robustness.
To reduce variance in generative outputs, each comparison is repeated three times and averaged. We validate this procedure by sampling 100 code pairs, performing manual robustness comparisons, and computing Cohen’s Kappa (Cohen, 1960), yielding 0.81, indicating near-perfect agreement. An alternative independent scoring approach (absolute scores 0–10 per snippet) gives a Kappa of 0.62, showing lower alignment with human judgment. All prompts, scoring instructions, and annotations are included in the released package (Rob, 2025). The evaluation uses GPT-4o via API calls, and details of hyperparameters (temperature, max tokens, etc.) are provided in the release. GPT-4o was chosen for its strong reasoning and code understanding capabilities, enabling scalable and reliable robustness assessment without model fine-tuning (Cheng et al., 2023).
Figure 2. Prompt for Evaluating Robustness
2.2.3. Statistical Metrics Analysis
Beyond LLM judgments, we employ statistical metrics to capture concrete aspects of robustness, specifically examining control expressions and exception handling as indicators of defensive programming practices.
Control Expressions Analysis. To assess the structure and complexity of control logic, we analyze atomic Boolean expressions, defined as the smallest evaluable Boolean conditions in control statements (e.g., if, for, while). Examples include simple checks like x > 5 or isValid(). Using abstract syntax trees (ASTs) extracted via Tree-sitter (Hendrik van Antwerpen, 2022), we identify all atomic Boolean expressions in both generated and reference code.
We define the following metric:
Average Atomic Boolean Expressions (AvgABE): measures the average number of atomic Boolean expressions per snippet:
@@ 2.2.3. Statistical Metrics Analysis
where $U_{i}$ is the set of atomic Boolean expressions in snippet $i$ and $N$ is the total number of snippets. A higher AvgABE indicates more explicit conditions, suggesting stronger robustness through enhanced boundary checks and input validation.
Exception Handling Analysis. Exception handling is a critical component of robust software. We define:
Exception Handling Adoption Rate (EHAR): the proportion of code snippets that include try-catch blocks:
@@
A higher EHAR reflects greater reliance on exception handling, which can help capture runtime errors but may also indicate heavier dependence on reactive rather than preventive mechanisms.
Table 2. Comparison of Robustness Metrics Across Models: “DSC” for DeepSeekCoder and “QWC” for Qwen2.5-Coder, “Gen” means Generated Code, “GT” means Ground True Code
Model EHAR AvgABE
@@
DSC-1.3B 0.02 0.03 1.39 2.04
DSC-6.7B 0.03 0.02 1.46 2.10
QWC-1.5B 0.01 0.03 1.62 2.12
QWC-7B 0.03 0.03 1.83 2.02
Figure 3. GPT-4o as a Judger: Results. “Generated Underperforms Human” indicates that the LLM judges human written code to be more robust than LLM generated code, the same applies to other labels.
@@ 2.2.4. Results
For the selected LLM-generated code, We examined the distribution of LLM Evaluator scores measuring the relative robustness of generated code compared to human written counterparts, and computed control expression and exception handling metrics relative to the reference code. The results are presented in Table 2 and Figure 3.
LLM-Based Automatic Evaluator Results Analysis. Figure 3 shows the distribution of tasks where generated code exhibits lower, equal, or better robustness compared to reference code. On average, 35.2% of generated code is less robust, 45.1% is equally robust, and 19.5% is more robust. The gap between cases where human code outperforms and underperforms generated code highlights substantial room for improvement.
For Qwen2.5-Coder-1.5B, 39.0% of compiled code is less robust than the reference code, while for Qwen2.5-Coder-7B, this proportion decreases to 28.5%. This suggests that increasing the parameter size of the Qwen2.5-Coder models leads to improved robustness in generated code. These results suggest that, overall, larger models tend to generate more robust code. When comparing models of similar parameter sizes, Qwen2.5-Coder exhibits superior robustness in code generation compared to DeepSeekCoder. Specifically, for Qwen2.5-Coder-1.5B, the proportion of code with “Human poorer than Generated” is 18.8%. In contrast, for DeepSeekCoder-1.3B, the proportion is 16.3%, respectively, indicating slightly weaker robustness. A similar trend is observed in the DeepSeekCoder series: for DeepSeekCoder 1.3B and 6.7B, the proportions of generated code with lower robustness compared to the reference code are approximately 41.8% and 31.2%, respectively.
Control Expression Analysis. We performed a Chi-square test (Pearson, 1900) to assess the association between the number of atomic expressions and code robustness. The result revealed a statistically significant association between the number of atomic expressions and the robustness of generated code (p ¡¡ 0.05). The corresponding Cramér’s V of 0.47 indicates a moderate-to-strong relationship, suggesting that a higher number of atomic expressions is meaningfully linked to increased robustness in generated code.
The AvgABE metric quantifies the average number of atomic Boolean expressions per code snippet. As indicated in Table 2, for the ground truth code, the AvgABE values across different subsets of model-generated code consistently remain around 2.0, suggesting that each ground truth snippet contains approximately two atomic Boolean expressions on average. In contrast, model-generated code tends to exhibit lower AvgABE values. Notably, DeepSeekCoder-1.3B demonstrates the lowest AvgABE value of 1.39. Among the evaluated models, Qwen2.5-Coder-7B performs relatively well, achieving an AvgABE of 1.83.
Additionally, we observe systematic differences across model series: the Qwen2.5-Coder series generally produces code with higher AvgABE values compared to the DeepSeekCoder Coder series. For example, Qwen2.5-Coder-1.5B generates code with an AvgABE value exceeding that of DeepSeekCoder-6.7B.
@@
Finding 1: LLM evaluator results show that LLM-generated code is on average 81.5% which robustness is less than or equal to human-written code. Robustness metrics further indicate that control expressions warrant more attention than exception handling, as they occur more frequently in real-world code and exhibit a larger performance gap.
@@ 2.3.1. Data Selection
The dataset used for this analysis originates from RQ1, consisting of 920 generated code snippets from four models (230 snippets per model). Given the high cost of full manual analysis, we use LLM evaluator as an initial filtering step to retain only code snippets exhibiting lower robustness than the reference implementation, resulting in a refined dataset with 364 code snippets for detailed manual inspection.
Further, to maintain the continuity and consistency of the analysis, we follow the task selection strategy employed in RQ1 2.2 and exclude all code snippets that cannot be successfully compiled. As a result, 220 code snippets are kept for further analysis.